OpenStudy (exoexoexo):

simplify (x^3y^4/xy^5)^3

OpenStudy (jhannybean):

Have you attempted the problem yet?

OpenStudy (exoexoexo):

my solution is y^3/x^6 but I don't know if it is correct or not

OpenStudy (jhannybean):

so your function is: \[\large \left(\frac{x^3y^4}{xy^5}\right)^3\] and you distribute the cubic power to all terms in the numerator and denominator : \[\large \left(\frac{x^6y^{12}}{x^3y^{15}}\right)\] now split up the x and y terms and simplify. \[\large \frac{x^6}{x^3} \cdot \frac{y^{12}}{y^{15}} = x^{6-3} \cdot y^{12-15} \] For negative powers, you stick them in the denominator.

OpenStudy (exoexoexo):

\[x ^{-9}\]

OpenStudy (exoexoexo):

(x^3y^4/xy^5)^-3